For a SDOF system with mass 1.4 kg, spring constant 4.2 N/m, and damping constant 1.5 N-sec/m, the
damped period is (two decimals, units).
Answer #1

natural circular frequency = (spring constant/mass) =3 = 1.732 rad/s

now critical damping constant = 2*mass*natural circular frequency = 2*1.4*1.732 = 4.8497 Nsec/m

damping factor = damping constant/critical damping constant = 1.5/4.8497 =0.3092

so damped circular frequency = natural circular frequency*(1-damping factor 2)

1.732 *(1-0.30922) =1.6471 rad/s

Time period of damping = 2*/damped circular frequency = 3.8146 seconds Ans